# Resource-Optimized Grouping Shadow for Efficient Energy Estimation

### v1

@@ A.3 Adaptive Resource-Optimized Grouping Shadow
2A fixed Pauli basis measurement recipe $\{\mathcal{P}_{\alpha}^{\otimes M_{\alpha}}\}_{1\leq\alpha\leq A}$ containing $\{X,Y,Z\}^{n\times M}$ for the classical shadow quantum measurement stage.
@@ Appendix C Fine Graining Shots Allocation
These results underscore the method’s robustness within a certain range of the parameter $\epsilon$. On the other hand, if one holds the product $\epsilon^{2}M$ fixed, the shots allocation ${w_{\alpha}}$ obtained by the convex optimization is fixed, and so is the confidence bound $\delta=\mathrm{Conf}\left[\epsilon_{0},{w_{\alpha}};M_{0},{\mathcal{H}}\right]$, which implies that this hyperparameter $\epsilon_{0}$ can be ascertained from a small number of measurements $M_{0}$. With $\epsilon_{0}^{2}M_{0}$ held fixed in  (1), the shot allocation $\{w_{\alpha}\}$ which minimizes the confidence bound is unchanged, meaning that if we increase the total number of shots to $M$ while setting $\epsilon=\epsilon_{0}\sqrt{M_{0}/M}$, the shots allocation $\{w^{\alpha}\}$ which minimizes the confidence bound $\delta=\mathrm{Conf}\left[\epsilon_{0},{w^{*}_{\alpha}};M_{0},{\mathcal{H}}\right]$ remains unchanged. Therefore, the algorithm can be improved by incorporating a coarse-grained search for the hyperparameter $\epsilon_{0}$ to find the minimum error of the energy estimation for a given state with a finite number of shot measurements $M_{0}$, where the energy estimation $\hat{\mathcal{H}}$ is obtained by quantum measurement using the measurement recipe $\{\mathcal{P}_{\alpha}^{\otimes M_{\alpha}^{0}}\}\in\{X,Y,Z\}^{n\times M_{0}},\leavevmode\nobreak\ (\sum M_{\alpha}^{0}=M)$, which is given by minimizing the confidence bound using a given $\epsilon_{0}$ following algorithm 3.3,
$\mathrm{Error}(\{\{\mathcal{P}^{(m)}_{\alpha}\}_{1\leq m\leq M^{0}_{\alpha}}\})=\sqrt{\mathbb{E}\left[\left(\hat{{\mathcal{H}}}(\{\mathcal{P}_{\alpha}^{\otimes M_{\alpha}^{0}}\})-\langle{\mathcal{H}}\rangle_{\rho}\right)^{2}\right]};$ (30)
@@
2A fixed Pauli basis measurement recipe $\{\mathcal{P}_{\alpha}^{\otimes M_{\alpha}}\}$ containing $\{X,Y,Z\}^{n\times M}$ for the classical shadow quantum measurement stage.
@@
8    $error[\epsilon_{cg}]=\mathrm{Error}(\{\mathcal{P}_{\alpha}^{\otimes M_{\alpha}^{cg}}\})$ (30)

### v2

@@ A.3 Adaptive Resource-Optimized Grouping Shadow
2A fixed Pauli basis measurement recipe $\{\mathcal{P}_{\alpha}^{\times M_{\alpha}}\}$ containing $\{X,Y,Z\}^{n\times M}$ for the classical shadow quantum measurement stage.
@@ Appendix C Fine Graining Shots Allocation
These results underscore the method’s robustness within a certain range of the parameter $\epsilon$. On the other hand, if one holds the product $\epsilon^{2}M$ fixed, the shots allocation ${w_{\alpha}}$ obtained by the convex optimization is fixed, and so is the confidence bound $\delta=\mathrm{Conf}\left[\epsilon_{0},{w_{\alpha}};M_{0},{\mathcal{H}}\right]$, which implies that this hyperparameter $\epsilon_{0}$ can be ascertained from a small number of measurements $M_{0}$. With $\epsilon_{0}^{2}M_{0}$ held fixed in  (1), the shot allocation $\{w_{\alpha}\}$ which minimizes the confidence bound is unchanged, meaning that if we increase the total number of shots to $M$ while setting $\epsilon=\epsilon_{0}\sqrt{M_{0}/M}$, the shots allocation $\{w^{\alpha}\}$ which minimizes the confidence bound $\delta=\mathrm{Conf}\left[\epsilon_{0},{w^{*}_{\alpha}};M_{0},{\mathcal{H}}\right]$ remains unchanged. Therefore, the algorithm can be improved by incorporating a coarse-grained search for the hyperparameter $\epsilon_{0}$ to find the minimum error of the energy estimation for a given state with a finite number of shot measurements $M_{0}$, where the energy estimation $\hat{\mathcal{H}}$ is obtained by quantum measurement using the measurement recipe $\{\mathcal{P}_{\alpha}^{\otimes M_{\alpha}}\}_{1\leq\alpha\leq A}\in\{X,Y,Z\}^{n\times M_{0}},\leavevmode\nobreak\ (\sum M_{\alpha}^{0}=M)$, which is given by minimizing the confidence bound using a given $\epsilon_{0}$ following algorithm 3.3,
$\mathrm{Error}(\{\mathcal{P}_{\alpha}^{\otimes M_{a}lp}\})=\sqrt{\mathbb{E}\left[\left(\hat{{\mathcal{H}}}(\{\mathcal{P}_{\alpha}^{\otimes M_{\alpha}}\}_{1\leq\alpha\leq A})-\langle{\mathcal{H}}\rangle_{\rho}\right)^{2}\right]};$ (30)
@@
2A fixed Pauli basis measurement recipe $\{\mathcal{P}_{\alpha}^{\times M_{\alpha}}\}$ containing $\{X,Y,Z\}^{n\times M}$ for the classical shadow quantum measurement stage.
@@
8    $error[\epsilon_{cg}]=\mathrm{Error}(\{\{\mathcal{P}^{(m)}_{\alpha}\}_{1\leq m\leq M^{cg}_{\alpha}}\})$ (30)